The membrane that surrounds a certain type of living cell has a surface area of 5.3 x 10^-9 m2 and a thickness of 1.5 x 10^-8 m. Assume that the membrane behaves like a parallel plate capacitor and has a dielectric constant of 5.3. (a) The potential on the outer surface of the membrane is +89.0 mV greater than that on the inside surface. How much charge resides on the outer surface? (b) If the charge in part (a) is due to K+ ions (charge +e), how many such ions are present on the outer surface?
1
Expert's answer
2020-09-10T11:58:50-0400

Concept

The capacitance of cell membrane is given by the expression:



"c=\t \\dfrac{k\\epsilon_{o}A }{d}"

The charge on the capacitor is given by:

"q= CV"

We have to charge on the outer surface of the membrane and also the number of ions present on it.


The Capacitance

"C=\t \\dfrac{k\\epsilon_{o}A }{d}"

"C=\t \\dfrac{(5.3)\\cdot(8.85*10^{-12})\\cdot(5.3*10^{-9})}{1.5*10^{-8}}"


"C=1.65*10^{-11}F"

Part(a):The charge on capacitor

"q=CV"


"q=(1.65*10^{-11}F)(89.0*10^{-3}V)"


"q=1.46*10^{-12}C"

Part(b): Number of Cations


From the law of quantization, we can find the number of cations:


"N= \dfrac{q}{e}"


"N= \dfrac{1.46*10^{-12}C}{1.6*10^{-19}C}"

"N=9.12*10^{6}"

Final answer:

The charge is "1.46*10^{-12}C" and number of cations are "9.12*10^{6}"